A steam boiler uses coal as a fuel in a furnace. The ultimate analysis (by mass) of the coal received is as follows:

C 80%, H2 2%, O2 4%, S 1.5%, ash 9% and moisture 3.5%.

Excess air supplied is 50%. Calculate the mass of air, and the mass of gas compositions in the flue gases per kg of coal burnt.
Answer #1

GIVEN:

Ultimate analysis of the coal received is tabulated below.

Table: 1

Component

Percentage by weight

C

80

H2

2

O2

4

S

1.5

Ash

9

Moisture

3.5

Excess air supplied = 50%

To FIND/TO PROVE:

1. The mass of air per kg coal burnt

2. The mass of gas compositions in the flue gases per kg coal burnt

SOLUTION:

STEP 1:

For complete combustion of 1 kg of coal, the amount of oxygen required and the products of combustion formed can be calculated as below.

Table: 2

Element Mass (kg.)

O2 Required (kg.)

Dry Products (kg.)

C = 0.8

0.8×83=2.13

0.8×113=2.93 CO2

H2 = 0.02

0.02×8=0.16

0.02×9=0.18 H2O

O2 = 0.04

-0.04

S = 0.015

0.015×1=0.015

0.015×2=0.03 SO2

                                                                                Total O2 = 2.265 kg

Considering 23% O2 present in the air.

Mass of air required for combustion=Total O20.23=2.2650.23=9.8478 kg.

Since 50% of excess air is supplied.

Therefore, Actual mass of air supplied can be calculated as follows

Actual mass of air=Mass of air required for combustion ×1.50                                                  =9.8478×1.5=14.7717 kg.

Therefore, mass of air per kg of coal burnt is 14.7717 kg

STEP 2:

Flue gases will be composed by N2, O2, CO2 and SO2

1. N2 Present in the products of combustion can be calculated as follows,

Considering 77% N2 present in the air.

Total mass of N2 in Dry products = N2 Present in air supplied.

Total mass of N2 in dry products= N2 Present in air supplied. =Actual mass of air ×0.77=14.7717 ×0.77=11.3742 kg

Therefore, mass of N2 per kg of coal burnt present in the flue gases is 11.3742 kg

2. O2 Present in the products of combustion can be calculated as follows,

Excess air supplied=Actual mass of air-mass of air required for combustion.

Excess air supplied=14.7717-9.8478=4.9239 kg

Considering 23% O2 Present in the air

O2 Present in the dry products= O2 present in the Excess air=Excess air supplied ×0.23=4.9239×0.23=1.1325 kg

Therefore, mass of O2 per kg of coal burnt present in the flue gases is 1.1325 kg

3. CO2 Present in the products of combustion is calculated in Table 2.

Therefore, CO2 Present in the products of combustion = 2.93kg

Therefore, mass of CO2 per kg of coal burnt present in the flue gases is 2.93 kg

4. SO2 Present in the products of combustion is calculated in Table 2.

Therefore, SO2 Present in the products of combustion = 0.03 kg

Therefore, mass of SO2 per kg of coal burnt present in the flue gases is 0.03 kg

FINAL ANSWER:

Mass of air per kg of coal burnt is 14.7717 kg

Mass of N2 per kg of coal burnt present in the flue gases is 11.3742 kg

Mass of O2 per kg of coal burnt present in the flue gases is 1.1325 kg

Mass of CO2 per kg of coal burnt present in the flue gases is 2.93 kg

Mass of SO2 per kg of coal burnt present in the flue gases is 0.03 kg